OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-7,45) and y-intercept (0,-200). Write your answer in this form: (x1,y1),(x2,y2). If necessary, round to the nearest hundredth. Use the method of completing the square to write the equation of the given parabola in this form: (y – k) = a(x – h)2 where a ≠ 0, (h,k) is the vertex, and x = h is the axis of symmetry.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (anonymous):

@phi @dan815 @IrishBoy123 @pooja195 @welshfella

OpenStudy (benlindquist):

find the common denominator

OpenStudy (benlindquist):

wrong person

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i just dont understand this

OpenStudy (phi):

did you learn the vertex form of an equation for a parabola?

OpenStudy (anonymous):

i have notes....let me see

OpenStudy (anonymous):

what it says

OpenStudy (anonymous):

well its asking for me to complete the square

OpenStudy (anonymous):

it confuses me though because they ask it in a different way

OpenStudy (anonymous):

@jabez177

jabez177 (jabez177):

@AlexandervonHumboldt2

OpenStudy (phi):

the form is \[ y= a(x-h)^2 + k\] where (h,k) is the vertex they tell you the vertex, so you can replace the h and k with numbers in that formula can you do that ?

OpenStudy (anonymous):

ehhh yeh

OpenStudy (anonymous):

there are 4 numbers though......

OpenStudy (anonymous):

i cant put both h and k in can i? i mean both h and k's i cant

OpenStudy (phi):

the vertex is (-7, 45) that means use h=-7 and k= 45 in the equation y = a(x-h)^2 + k so replace the h and k letters with numbers. what do you get ?

OpenStudy (anonymous):

ohh okay let me try

OpenStudy (anonymous):

y=a(x-7)^2+45?

OpenStudy (phi):

it should be x - h in this case x- (-7)

OpenStudy (anonymous):

im really lost like this is algebra 2 and up. i only know algebra 1 or less.

OpenStudy (phi):

can you fix your equation? it is not x-7

OpenStudy (anonymous):

ehh y=a(-7)+45?

OpenStudy (phi):

no, don't get lost. you want a(x-h)^2 + 45 h is -7 can you simplify x - -7 ? minus a minus is plus, remember?

OpenStudy (anonymous):

yeah but im not a very mathematical person. all i see are letters ,numbers,and signs

OpenStudy (anonymous):

and then my brain goes in 5 different ways

OpenStudy (phi):

examples: 1 - 1 = 0 1- -1 = 1+1= 2

OpenStudy (anonymous):

ohh

OpenStudy (phi):

x- -7 is x+7

OpenStudy (phi):

so far you have \[ y= a(x+7)^2 + 45\] they also tell you (0,-200) is on the curve "plug in" x=0 and y= -200 into your equation. what do you get ?

OpenStudy (anonymous):

okay so let me do this

OpenStudy (anonymous):

im working it out on paper

OpenStudy (anonymous):

uhm is it -200=a(0+7)^2+45?

OpenStudy (phi):

yes. now remember order of operations? first do the parens: 0+7

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

do i square the -7?

OpenStudy (anonymous):

so that would be -200=a(-49)+45?

OpenStudy (phi):

you get -200= a(7)^2 + 45 (it's +7 or just 7) but yes, now do 7*7

OpenStudy (anonymous):

which is 49

OpenStudy (phi):

yes uhm is it -200=a(0+7)^2+45? becomes -200 = 49a + 45 now add -45 to both sides

OpenStudy (phi):

like this -200 + -45 = 49a + 45 + -45

OpenStudy (anonymous):

oooh yeah okay hold on

OpenStudy (anonymous):

is it a=-200/49?

OpenStudy (phi):

go slower. you have the right idea, but you have to be careful -200 + -45 = 49a + 45 + -45 on the left side, what is -200 + -45

OpenStudy (anonymous):

-245

OpenStudy (phi):

and on the right side 45 + -45 is 0 so you have -245 = 49a + 0 adding 0 does not do anything so it's just -245= 49a or 49a = -245

OpenStudy (phi):

now you can divide both sides by 49

OpenStudy (anonymous):

so divide -245 by 49 and 49 by 49 which will be 1/5

OpenStudy (anonymous):

i think...

OpenStudy (phi):

you do \[ \frac{49}{49} a = \frac{-245}{49} \]

OpenStudy (anonymous):

yeah i did that so a=5?

OpenStudy (phi):

you mean a= -5 , right ?

OpenStudy (anonymous):

oh yeah lol

OpenStudy (anonymous):

okay so i have to solve for x and y?

OpenStudy (anonymous):

and one point is -5

OpenStudy (phi):

you found a= -5 so the equation of the curve \[ y= a(x+7)^2 + 45 \] is \[ y = -5(x+7)^2 +45\] this equation tells us everything we want to know about the parabola.

OpenStudy (anonymous):

oooh so i just plug it in now?

OpenStudy (phi):

They want Find the x-intercepts of the parabola the x-intercepts are the x values when y is 0 in other words, we have to "solve for x" when y is 0: \[ y = -5(x+7)^2 +45\\ -5(x+7)^2 +45=0\] I would first add -45 to both sides

OpenStudy (anonymous):

it depens what we are looking for like is it x

OpenStudy (anonymous):

okay so its -7 then

OpenStudy (phi):

no, when x is 0, we get the y-intercept. this is harder. we set the y=0 and we have to find the x that makes y zero. we start with \[ -5(x+7)^2 +45 = 0\] now add -45 to both sides

OpenStudy (anonymous):

okay well -5+ -45 is 40

OpenStudy (anonymous):

oops i mean -50

OpenStudy (phi):

write down the entire equation (including the x stuff!) \[ -5(x+7)^2 +45 + -45= 0+-45\]

OpenStudy (anonymous):

okay

OpenStudy (phi):

to review, we start with \[ -5(x+7)^2 +45 = y\] and we set y =0 (this is how to find the x-intercepts) \[ -5(x+7)^2 +45 = 0\] now we add -45 to both sides \[ -5(x+7)^2 = -45 \] next , divide both sides by -5

OpenStudy (anonymous):

okay i got 1 and 9

OpenStudy (phi):

you get \[ (x+7)^2 = \frac{-45}{-5} \]

OpenStudy (phi):

or \[ (x+7)^2 = 9\] now take the square root of both sides

OpenStudy (anonymous):

square root of both sides?

OpenStudy (phi):

yes. do you know how to do that ?

OpenStudy (anonymous):

just 7^2 or 9^2

OpenStudy (anonymous):

or what goes into both evenly

OpenStudy (anonymous):

ugh i give up

OpenStudy (phi):

square root "undoes" the square in other words you do \[ \sqrt{(x+7)^2} = \sqrt{9} \]

OpenStudy (phi):

and it's worth remembering \[ \sqrt{x^2} = x \] or in this case \[ \sqrt{(x+7)^2}= x+7\]

OpenStudy (phi):

for the square root of 9, what number times itself is 9 ? 3*3 is 9 that means 3 is the square root. or if you are confused, you can use a calculator you get \[ x+7= \pm 3 \] we usually show both -3 and +3 work: -3*-3 is 9 and so is 3*3 = 9

OpenStudy (phi):

we have two different equations: x+7 = -3 and x+7 = 3 solve each one. can you do that ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

hold on